OpenStudy (anonymous):

Solve the System 3x-y=9 2y+z=8 x+4z=12

OpenStudy (anonymous):

1 0 4 12 0 2 1 8 3 -1 0 9

OpenStudy (anonymous):

subtract 3*row 1 from row 3

OpenStudy (anonymous):

1 0 4 12 0 2 1 8 0 -1 -12 -27

OpenStudy (anonymous):

add 2*row 3 to row 2

OpenStudy (anonymous):

1 0 4 12 0 0 -23 -46 0 -1 -12 -27

OpenStudy (anonymous):

z = -46/-23 = 2

OpenStudy (anonymous):

x + 4z = 12 x = 12 - 4(2) = 4

OpenStudy (anonymous):

-y -12z = -27 y = 27 - 12(2) = 3

OpenStudy (anonymous):

x = 4 y = 3 z = 2